Grade 11Mechanics

If a vector A= cos Omega t I + sin Omega t j and B = cosomega t/2+I + sin Omega t /2 are function of time then the value of the at which they are perpendicular to each other

Payal Pradhan
8 Years agoGrade 11

1 Answer

Mahesh Koilada
8 Years ago
A and B are perpendicular at time t = T (say). Then dot product of vectors, A.B=0.
 
So, (cos 2x) * (cos x) + (sin 2x) * (sin x) = 0   { here x = omega T /2 (say) }
 
simplifying above equation, we get
 
-1/( tan x) = tan 2x = 2 tan x / (1- (tan x) ^ 2) and this gives, on simplification, the following
 
(tan x)^2 = -1. This implies tan x is imaginary. Therefore (tan omega T /2) is imaginary.
 
But we learned that “tan” is a real function. That mens, in this case, ‘T’ is not defined.
 
Hence vectors A & B can never be perpendicular to each other at any time ‘t’.
 
All the best.